# On two maximally entangled couples

### v1

@@ Abstract.
F.H.’s research was funded in whole or in part by the National Science Centre, Poland 2024/54/E/ST2/00451. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. J.S. was supported by Grant PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by ”ERDF A way of making Europe” as well as by the Basque Government through Grant No. IT1470-22.
@@ Introduction
The existence of absolutely maximally entangled (AME) states (and in particular that of four-qubit AME states) is, in a sense, one of the simplest instances of the quantum marginal problem [Kly06] and beautifully illustrates a key difficulty at the core of all of many-body physics: It asks whether there exists a pure $n$-qubit state, such that every bipartition is maximally entangled. Naturally, in the case of two and three qubits, the Bell and Greenberger-Horne-Zeilinger states satisfy this requirement. But what about four qubits?
In a seminal article (revisited in Proof 1: Higuchi-Sudbery), Higuchi and Sudbery found that no such state can possibly exist. This opened up a field of research that dealt with both bounds and constructions on absolutely maximally entangled and $k$-uniform states, the latter having all $k$-body marginals maximally mixed [GAL+15, HGS17, RBB+22]. If a whole subspace satisfies such condition, we are dealing with a pure quantum code of distance $k+1$ [HG20]. Indeed it was early recognized by Scott that methods of coding theory are useful in the characterization of quantum entanglement [Sco04]. We present here two proofs inspired by coding theory. Proof 5: Shadow inequalities relies on Rains’ shadow inequalities [Rai99], a monogamy relation constraining the purities in multipartite quantum states. Proof 7: Lovász bound rests on the Lovász theta number, a quantity traditionally used to bound the independence number and thus the capacity of a graph, but which has recently shown to be also applicable to quantum codes [MNH24].
@@
While this collection of proofs shows that a lot is known, it is still difficult to pinpoint the essence of this non-existence fact. Although we are perfectly able to outline the problem and can elucidate its technical solution by connecting it to many different facts and mathematical properties of quantum many-body systems, we lack a simple intuition for how local features of the parts relate to (and constrain) the global structure of the system as a whole.
@@
We review the following proofs. To our knowledge, Proof 3: Opposite two-body correlations and Proof 4: Polynomial Invariants of Four Qubits are presented here for the first time.
@@
In their remarkable work, Higuchi and Sudbery [HS00] analyzed the possibility of maximum entanglement among four qubits. While entanglement for two and three qubits was well understood at the time, for $n>3$ qubits not much was known. They recognized that the obvious $n$-qubit entangled Greenberger-Horne-Zeilinger (GHZ) state is probably not the ’most entangled’, where that concept had yet to be specified. They did the latter by requiring the maximum number of maximally mixed reductions – from then on this has been the common definition for absolute maximal entanglement (AME) [HC13]. They noticed that it is not obvious that such a state would exist: The Bell state ($n=2$) and the GHZ state ($n=3$) correspond to the definition. For four qubits there at least exists the state
$\mathinner{|\Psi\rangle}\ =\ \frac{1}{2}\left(\mathinner{|0000\rangle}+\mathinner{|0011\rangle}+\mathinner{|1001\rangle}+\mathinner{|1110\rangle}\right)\ \ ,$
@@
(2) $a_{jklm}\ =\ \frac{1}{2}\left(U_{1}\right)_{jl,km}\ =\ \frac{1}{2}\left(U_{2}\right)_{jk,lm}\ =\ \frac{1}{2}\left(U_{3}\right)_{jm,kl}\ \ ,$
where $U_{1}$, $U_{2}$ and $U_{3}$ are unitary matrices. By applying local unitaries to $\mathinner{|\psi\rangle}$ it can be achieved that
(3) $a_{1000}\ =\ a_{0100}\ =\ a_{0010}\ =\ a_{0001}\ \ .$
Subsequently they argue by using the orthogonality conditions of the columns and rows of the unitary matrices $U_{j}$ ($j=1,2,3$) that more coefficients of $\mathinner{|\psi\rangle}$ must be zero. With the remaining non-zero elements it is impossible to simultaneously guarantee unitarity of $U_{1}$, $U_{2}$ and $U_{3}$, which concludes their proof.
@@
The proof follows the strategy of: 1) recognizing that all marginals of an AME state are proportional to projectors, 2) expanding the density matrix into correlations of different weights, and 3) splitting the resulting equation by an even-odd parity rule on the weights of anti-commutators of Pauli matrices.
@@
Suppose there exists a a four-qubit AME state $\mathinner{|\phi\rangle}$. Let $\varrho_{123}=\operatorname{tr}_{4}(|\phi\rangle\langle\phi|)$ be its three-body reduced density matrix on the first three systems. It can be written as
@@
where $P_{3}$ contains three-body terms only. That is, it has the form
(5) $P_{3}=\sum_{\alpha,\,\beta,\,\gamma\neq 0}\,c_{\alpha\beta\gamma}\,\sigma_{\alpha}\otimes\sigma_{\beta}\otimes\sigma_{\gamma}\,.$
where the sum is over all Pauli matrices $\sigma_{\alpha},\sigma_{\beta},\sigma_{\gamma}\neq\mathds{1}$.
@@
[…]
@@
[…]
Now the second equation yields clearly a contradiction: necessarily $P_{123}\neq 0$ due to the fact that $\varrho_{123}$ cannot have full rank, which follows from the Schmidt decomposition of a pure state. As a consequence, a four-qubit AME state does not exist. ∎
@@
The even-odd Lemma is then used to split two projector relations as in Eq. (6) each into two contributions, for which the eigenvector relations provide a contradiction.
@@
For this proof we need a short lemma.
@@
We return to the non-existence of a four-qubit AME state.
@@
A four-qubit AME state corresponds to a pure quantum error-correcting code with parameters $(\!(4,1,3)\!)_{2}$[CRSS98, Sco04]. Any pure code with parameters $(\!(n,K,d)\!)_{D}$ can be propagated by a partial trace to a pure $(\!(n-1,DK,d-1)\!)_{D}$ code [Rai98, Theorem 19]. Propagating the $(\!(4,1,3)\!)_{2}$ by a partial trace on the last party yields a pure code with parameters $(\!(3,2,2)\!)_{2}$, represented by the associated density matrix $\varrho_{ABC}$. By the Schmidt decomposition, this reduced state has eigenvalues $(\tfrac{1}{2},\tfrac{1}{2})$ and thus fulfills the relation $\varrho_{ABC}^{2}=\frac{1}{2}\varrho_{ABC}$. We can write this state as a convex combination,
@@
Now note that because $\varrho_{ABC}=(|v\rangle\langle v|+|w\rangle\langle w|)/2$ is two-uniform (being the reduction of a two-uniform state), we must necessarily have that $Q_{2}=-P_{2}$ in Eq. (Proof.). Thus for $|w\rangle\langle w|$, we need to make the Ansatz
@@
By Eq. (Proof.), $Q_{3}$ has the non-zero eigenvalues $\{4,-4\}$. Now recall that that $\lambda_{\max}(A+B)\leq\lambda_{\max}(A)+\lambda_{\max}(B)$ for all hermitian matrices $A,B$. Then $\mathinner{|w\rangle}$ being pure requires that $\lambda_{\text{max}}(Q_{3})\geq 6$ so that $\lambda_{\text{max}}(\mathds{1}-P_{2}+Q_{3})=8$. This leads to a contradiction, as by Eq. (Proof.) one has $\lambda_{\text{max}}(Q_{3})=4$. This ends the proof.
@@
Polynomial invariants are relevant for the quantification of entanglement because they allow for a straightforward construction of measures for different classes of genuine multipartite entanglement [VDDM03, ES14]. For states of four qubits it is possible to construct the full generating set of invariants under the group SL$(2,C)^{\otimes 4}$ as was demonstrated by Luque and Thibon [LT03]. It consists of four elements, one degree-2 invariant, two degree-4 invariants and one degree-6 invariant. By writing a four-qubit state in the computational basis, $\mathinner{|\psi\rangle}=\sum a_{j_{1}j_{2}j_{3}j_{4}}\mathinner{|j_{1}j_{2}j_{3}j_{4}\rangle}$ (where $j_{k}\in\{0,1\}$) one obtains three degree-4 invariants [LT03]
@@
(22a) $\rho_{12}\ =$ $\mathsf{L}^{T}\mathsf{L}^{*}\ \ ,$
(22b) $\rho_{24}\ =$ $\mathsf{M}\mathsf{M}^{\dagger}\ \ ,$
(22c) $\rho_{14}\ =$ $\mathsf{N}^{T}\mathsf{N}^{*}\ \ .$
For an AME state we must have $\rho_{12}=\rho_{24}=\rho_{14}=\frac{1}{4}\mathds{1}_{4}$. These equalities imply that the matrices $\mathsf{L}$, $\mathsf{M}$, and $\mathsf{N}$, up to a factor $\frac{1}{2}$, must be unitary. Accordingly, the columns and rows of each matrix form an orthogonal basis, with all vectors normalized to $\frac{1}{2}$. Moreover, we can choose the global phase of $\mathinner{|\psi\rangle}$ so that, because of Eq. (21),
@@
[…]
Assume for the moment that $a_{0011}\neq 0$ (we discuss the opposite case below). The orthogonality conditions then require the following structure of $\mathsf{L}$ (up to sign changes and column or row permutations),
@@
This proof uses the shadow inequality by Rains [Rai99, Rai00]: Denote by $\varrho_{S}=\operatorname{tr}_{S^{c}}(\varrho)$ the marginal of $\varrho$ on subsystem $S$. The inequality states that for all states $\varrho$ and subsets $T\subseteq\{1,\dots n\}$, it holds that 111 This trace inequality has been promoted to an operator inequality in Ref. [EHGS18].
@@
The non-negativity of this expression can be understood by projecting $\varrho^{\otimes 2}$ locally onto symmetric and anti-symmetric subspaces given by $\tfrac{1}{2}(\mathds{1}+\operatorname{swap})$ and $\tfrac{1}{2}(\mathds{1}-\operatorname{swap})$ and taking the trace [Rai98].
@@
[…]
@@
Gour and Wallach observed that for a four-qubit state [GBS07],
@@
where $\tau_{ABCD}=|\!\mathinner{\langle\psi|}\sigma_{y}^{\otimes 4}\mathinner{|\psi^{*}\rangle}\!|^{2}$. However, a state with maximally mixed two-qubit marginals must have $\tau_{1}=1$ and $\tau_{2}=3/2$, which is in contradiction to inequality (34). ∎
@@
This proof follows the strategy of Ref. [MNH24, Corollary 9], which provides a complete semidefinite programming hierarchy for the existence of qubit quantum codes with given parameters $(\!(n,K,d)\!)$. Here, $n$ denotes the number of qubits the code contains, $K$ is the dimension of the code space, and $d$ the distance. Here, we only need a relaxation of this hierarchy, which is given by the Lovász theta number.
Let $G$ be a graph of $N$ vertices, where we write $i\sim j$ if $\{i,j\}$ is an edge in the graph. The Lovász theta number is given by the following semidefinite program [GL17]:
@@
It is known that $\vartheta$ bounds the independence number of a graph, that is the maximum size of a set in which no two vertices are connected, as $\alpha(G)\leq\vartheta(G)$.
@@
Our aim is to construct a anti-commutativity graph $G$ that captures the correlations of the AME state, and to show that its theta number is too small for it to correspond to a pure state. Denote by $\sigma_{x},\sigma_{y},\sigma_{z}$ the three Pauli matrices, and by $\sigma_{0}$ the identity matrix. Let $\mathcal{P}_{n}$ be the $n$-qubit Pauli basis, that is the set of the $4^{n}$ elements of the form $E_{\alpha}=\sigma_{\alpha_{1}}\otimes\dots\otimes\sigma_{\alpha_{n}}$, where $\sigma_{\alpha_{i}}\in\{\sigma_{0},\sigma_{x},\sigma_{y},\sigma_{z}\}$. With some abuse of notation, map the multi-index $\alpha$ to an integer between 00 and $14^{n}-$ 1, so that $E_{0}=\mathds{1}$.
@@
where $10\leq\alpha,\beta\leq 4^{4}-$ 1. It can be seen that $\Gamma$ is positive semidefinite, which can be seen by expanding
@@
due to the fact that the Pauli operators are unitary and hermitian, so that $E_{\alpha}^{2}=\mathds{1}$ 222 Note that all Pauli operators are hermitian so we can safely drop any daggers.. Therefore the matrix $\Gamma$ satisfies: $\Gamma_{00}=1$ and $\Gamma_{\alpha 0}=\Gamma_{\alpha\alpha}$ for all $\alpha$. Furthermore, from the normalization $\operatorname{tr}(\varrho^{2})=1$ it follows that
@@
$\Gamma_{\alpha\beta}$ $=0\quad\quad\text{if}\quad 0<\operatorname{wt}(E_{\alpha}E_{\beta})<d\,,\quad$
(40) $\Gamma_{\alpha\alpha}$ $=0\quad\quad\text{if}\quad 0<\operatorname{wt}(E_{\alpha})<d\,.$
We observe that the second condition also implies that if $0<\operatorname{wt}(E_{\alpha})<d$ then also $\Gamma_{\alpha\beta}$ for all $\beta$. This is due to the fact that if a positive semidefinite matrix has a diagonal entry equal to zero, then all entries in the same row or column are zero too. Note that one could also obtain this constraint directly from the definition of $\Gamma_{\alpha\beta}$.
@@
In light of the conditions derived in Eqs. (Proof.) and (41), let us write
@@
Then if a four qubit AME state exists, the following semidefinite program must, due to Eq. (39), reach a value of at least $12^{4}-$ 1:
@@
But Eq. (Proof.) is nothing else than the Lovász theta number [Eq. (Proof 7: Lovász bound)] for the anti-commutativity graph given by the $n$-qubit Pauli basis with the identity matrix removed, $\mathcal{P}_{n}\backslash\{\mathds{1}\}$, and additional edges among low-weight products. Consequently, for an AME state to exist, we need to have $\vartheta(G)+1\geq 16$. However, for the anti-commutativity graph of the four-qubit Pauli group $\mathcal{P}_{4}\backslash\{\mathds{1}\}$, $\vartheta(G)+1=8$. As a consequence, a four-qubit AME state does not exist. ∎

### v2

@@ Abstract.
F.H.’s research was funded in whole or in part by the National Science Centre, Poland 2024/54/E/ST2/00451 and by the Polish National Agency for Academic Exchange under the Strategic Partnership Programme grant BNI/PST/2023/1/00013/U/00001. For the purpose of Open Access, the author has applied a $C$-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. J.S. was supported by Grant PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by ”ERDF A way of making Europe” as well as by the Basque Government through Grant No. IT1470-22. The corresponding author is FH
@@ Introduction
The existence of absolutely maximally entangled (AME) states, and in particular that of four-qubit AME states, is one of the simplest instances of the quantum marginal problem [Kly06] and beautifully illustrates a key difficulty at the core of all of many-body physics: It asks whether there exists a pure $n$-qubit state (more generally, $n$-qudit state), such that every bipartition is maximally entangled; or equivalently, for a pure $n$-qubit state such that all reduced density matrices  [^1] Here we use the words reduced density matrices and marginals interchangeably. A $k$-body reduced density matrix or $k$-body marginal is obtained after partially tracing out all but $k$ subsystems. of size $\lfloor\tfrac{n}{2}\rfloor$ are maximally mixed. Naturally, in the case of two and three qubits, the Bell and Greenberger-Horne-Zeilinger states satisfy this requirement. But what about four qubits?
In a seminal article (revisited in Proof 1: Higuchi-Sudbery), Higuchi and Sudbery found that no such state can possibly exist. This opened up a field of research that dealt with both bounds and constructions on absolutely maximally entangled and $k$-uniform states, the latter having all $k$-body marginals maximally mixed [GAL+15, HGS17, RBB+22]. We point to the review [RMBR+25] for a comprehensive background on absolutely maximally entangled states, in particular regarding constructions.
Let us sketch connections of this problem to quantum coding theory: if every vector in a subspace has maximally mixed $k$-body marginals, we are dealing with a pure quantum code of distance $k+1$ [HG20]. Indeed it was early recognized by Scott that methods of coding theory are useful in the characterization of quantum entanglement [Sco04]. We present here two proofs inspired by coding theory. Proof 5: Shadow inequalities relies on Rains’ shadow inequalities [Rai99], a monogamy relation constraining the purities in multipartite quantum states. Proof 7: Lovász bound rests on the Lovász theta number, a quantity traditionally used to bound the independence number and thus the capacity of a graph, but which has recently shown to be also applicable to quantum codes [MNH24]. This is a quite unexpected appearance of an invariant from graph theory, which here is used for a problem of continuous character, i.e. for the non-existence of a rank-$1$ subspace with some given properties.
@@
These different approaches highlight that there often exists more structure in these problems than strictly necessary for a proof. Furthermore, we see that entanglement theory, and in particular absolutely maximally entangled states, manages to tie together very different subfield of mathematics. While this collection of proofs shows that a lot is known, it is still difficult to pinpoint the essence of this non-existence fact. Although we are perfectly able to outline the problem and can elucidate its technical solution by connecting it to many different facts and mathematical properties of quantum many-body systems, we lack a simple intuition for how local features of the parts relate to (and constrain) the global structure of the system as a whole.
@@
In the remainder of the article we review seven proofs for this fact. To our knowledge, Proof 3: Opposite two-body correlations and Proof 4: Polynomial Invariants of Four Qubits are presented here for the first time.
@@
In their remarkable work, Higuchi and Sudbery [HS00] analyzed the possibility of maximum entanglement among four qubits. While entanglement for two and three qubits was well understood at the time, for $n>3$ qubits not much was known. They recognized that the obvious $n$-qubit entangled Greenberger-Horne-Zeilinger (GHZ) state is probably not the “most entangled”, where that concept had yet to be specified. They did the latter by requiring the maximum number of maximally mixed reductions – from then on this has been the common definition for absolute maximal entanglement (AME) [HC13]. They noticed that it is not obvious that such a state would exist: The Bell state ($n=2$) and the GHZ state ($n=3$) correspond to the definition. For four qubits there at least exists the state
$\mathinner{|\Psi\rangle}\ =\ \frac{1}{2}\left(\mathinner{|0000\rangle}+\mathinner{|0111\rangle}+\mathinner{|1001\rangle}+\mathinner{|1110\rangle}\right)\ \ ,$
@@
(2) $a_{jklm}\ =\ \frac{1}{2}\left(U_{1}\right)_{jk,lm}\ =\ \frac{1}{2}\left(U_{2}\right)_{jl,km}\ =\ \frac{1}{2}\left(U_{3}\right)_{jm,kl}\ \ ,$
where $U_{1}$, $U_{2}$ and $U_{3}$ are two-qubit unitary matrices. By applying local unitaries to $\mathinner{|\psi\rangle}$ it can be achieved that
(3) a1000=a0100= $a_{1000}\ =\ a_{0100}\ =\ 0\$=\ $a_{0001}\ \$.
To see that this is possible consider only the first equality, $a_{1000}=a_{0100}=0$. The four components $a_{mn00}$, $m,n\in\{0,1\}$, may formally be viewed as a two-qubit vector. As is well known the Schmidt form of this vector can be obtained by applying local unitaries to the qubits, corresponding here to unitary operations on the first and second qubit of $\mathinner{|\psi\rangle}$, which yield the first equality. Here we need to make sure that $|a_{0000}|$ increases, that is, we choose the local unitaries such that the new $a_{0000}$ equals the larger Schmidt coefficient. An analogous argument holds for the second part of the equality treated separately (by considering $a_{00pq}$) and is achieved by applying local unitaries to the third and fourth qubit of $\mathinner{|\psi\rangle}$. Again we choose the unitaries such that $|a_{0000}|$ increases. By iterating this procedure, one achieves the maximal $a_{0000}$, at which point both parts of the equality hold simultaneously.
Subsequently Higuchi and Sudbery argue by using the orthogonality conditions of the columns and rows of the unitary matrices $U_{j}$ ($j=1,2,3$) that more coefficients of $\mathinner{|\psi\rangle}$ must be zero. With the remaining non-zero elements it is impossible to simultaneously guarantee unitarity of $U_{1}$, $U_{2}$ and $U_{3}$, which concludes their proof.
@@
The proof follows the strategy of: 1) recognizing that all marginals of an AME state are proportional to projectors, and thus satisfy a quadratic equation; 2) expanding the density matrix into correlations of different weights; and 3) splitting the quadratic equation by an even-odd parity rule on the weights of anti-commutators of Pauli matrices leading to a contradiction.
@@
Suppose there exists a a four-qubit AME state $\mathinner{|\phi\rangle}$. Let $\varrho_{123}=\operatorname{tr}_{4}(|\phi\rangle\langle\phi|)$ be its three-body reduced density matrix on the first three systems. Since the state $\mathinner{|\phi\rangle}$ is AME, all of its two-body reduced states are maximally mixed, $\operatorname{tr}_{1}(\varrho_{123})=\operatorname{tr}_{2}(\varrho_{123})=\operatorname{tr}_{3}(\varrho_{123})=(\mathds{1}\otimes\mathds{1})/4$. Therefore $\varrho_{123}$ can be written as
@@
where $P_{3}$ contains three-body terms only. That is, $P_{3}$ contains only tensor products of three non-identity Pauli matrices $\sigma_{\alpha},\sigma_{\beta},\sigma_{\gamma}\neq\mathds{1}$,
(5) $P_{3}=\sum_{\alpha,\,\beta,\,\gamma\neq 0}\,c_{\alpha\beta\gamma}\,\sigma_{\alpha}\otimes\sigma_{\beta}\otimes\sigma_{\gamma}\,$,
so that $P_{3}$ vanishes under any partial trace applied to any single subsystem.
@@
$\{\text{even},\text{even}\}$ $\rightarrow\text{even}\,$,
$\{\text{odd},\text{odd}\}$ $\rightarrow\text{even}\,$,
(9) $\{\text{even},\text{odd}\}$ $\rightarrow\text{odd}\,.$
@@
(10) $P_{3}^{2}=\frac{1}{2}\{P_{3},P_{3}\}$ $=3\mathds{1}\quad\quad\text{(even)}$
(11) $0$ $=P_{3}\quad\quad\text{(odd)}$
Now the second equation yields clearly a contradiction: necessarily $P_{3}\neq 0$ due to the fact that $\varrho_{123}$ cannot have full rank, which follows from the Schmidt decomposition of a pure state. As a consequence, a four-qubit AME state does not exist. ∎
@@
where $S^{c}$ denotes the complement of $S$ in $\{1,\dots,n\}$. The even-odd Lemma is then used to split two projector relations as in Eq. (6) each into two contributions, for which the eigenvector relations provide a contradiction.
@@
We need a short lemma, originally proven in Ref. [SM96], which we now derive through an iterated Schmidt decomposition.
@@
We recall the definition of quantum error-correcting codes. A code with parameters $(\!(n,K,\delta)\!)_{2}$ is able to encode a $K$-dimensional system into $n$ qubits, such that all errors acting on at most $\lfloor\tfrac{\delta-1}{2}\rfloor$ subsystems can be corrected. Then $n$ is block-length, $K$ the dimension, and $\delta$ the distance of the code. More precisely, a subspace of $(\mathds{C}^{2})^{\otimes n}$ is a $(\!(n,K,\delta)\!)_{2}$ code, if and only if the Knill-Laflamme conditions hold for all Pauli strings $E$ with weight $\text{wt}(E)<\delta$,
(13) $\Pi E\Pi=c_{E}\Pi\,,\quad\quad\quad c_{E}\in\mathds{C}\,$,
where $\Pi$ is the projector associated with the code subspace. Here $\operatorname{wt}(E)$ is the number of subsystems a tensor-product operator $E$ acts non-trivially on.
A code is said to be pure if it additionally satisfies $c_{E}=\operatorname{tr}(E)/2^{n}$. Thus for a pure code, $\operatorname{tr}_{S^{c}}(\Pi)$ is maximally mixed for all $S\subseteq\{1,\dots,n\}$ of size $|S|<\delta$ and where $S^{c}$ is the complement of $S$ in $\{1,\dots,n\}$. This can be seen by writing $E=E_{S}\otimes\mathds{1}_{S^{c}}$ and taking the trace over Eq. (13),
(14) $\operatorname{tr}[\Pi E\Pi]=\operatorname{tr}[\Pi(E_{S}\otimes\mathds{1}_{S^{c}})]=\operatorname{tr}[\operatorname{tr}_{S^{c}}(\Pi)E_{S}]=\operatorname{tr}(E)/2^{n}\operatorname{tr}(\Pi)\,$,
which vanishes for all $E\neq\mathds{1}$. Thus $\operatorname{tr}_{S^{c}}(\Pi)$ is maximally mixed for all $|S|<\delta$.
It is known that if there exists a pure code with parameters $(\!(n,K,d)\!)_{D}$, then there also exists a pure $(\!(n-1,DK,d-1)\!)_{D}$ code [Rai98, Theorem 19]. This is known as code propagation, and the new code (sub-)space is obtained by taking a partial trace over a single subsystem of the old code space. After the partial trace only n1n-1 subsystems remain and thus the new block-length is n1n-1, while it can be shown that this procedure also decreases the distance by one with the dimension of the code space increasing by a factor of the local dimension that has been traced out.
We now show the non-existence of four-qubit AME states.
@@
Having all two-body marginals maximally mixed, a four-qubit AME state corresponds to a pure quantum error-correcting code with parameters $(\!(4,1,3)\!)_{2}$[CRSS98, Sco04]. Propagating the $(\!(4,1,3)\!)_{2}$ by a partial trace on the last party yields a pure code with parameters $(\!(3,2,2)\!)_{2}$, represented by the associated density matrix $\varrho_{ABC}$. By the Schmidt decomposition, this reduced state has eigenvalues $(\tfrac{1}{2},\tfrac{1}{2})$ and thus fulfills the relation $\varrho_{ABC}^{2}=\frac{1}{2}\varrho_{ABC}$. We can write this state as a convex combination,
@@
Now note that because $\varrho_{ABC}=(|v\rangle\langle v|+|w\rangle\langle w|)/2$ is two-uniform (being the reduction of a two-uniform state), we must necessarily have that $Q_{2}=-P_{2}$ in Eq. (Proof 3: Opposite two-body correlations). Thus for $|w\rangle\langle w|$, we need to make the Ansatz
@@
By Eq. (Proof 3: Opposite two-body correlations), $Q_{3}$ has the non-zero eigenvalues $\{4,-4\}$. Now recall that that $\lambda_{\max}(A+B)\leq\lambda_{\max}(A)+\lambda_{\max}(B)$ for all hermitian matrices $A,B$. Then $\mathinner{|w\rangle}$ being pure requires that $\lambda_{\text{max}}(Q_{3})\geq 6$ so that $\lambda_{\text{max}}(\mathds{1}-P_{2}+Q_{3})=8$. This leads to a contradiction, as by Eq. (Proof 3: Opposite two-body correlations) one has $\lambda_{\text{max}}(Q_{3})=4$. This ends the proof. ∎
@@
Polynomial invariants are polynomial expressions of the coefficients of a pure multi-party quantum state (with local dimension $d$), which remain invariant under the application of local SL($d$,$C$) operations. These operations are represented by complex $d\times d$ matrices with determinant equal to 1. Polynomial invariants are relevant for the quantification of entanglement because they allow for a straightforward construction of measures for different classes of genuine multipartite entanglement [VDDM03, ES14].
[…]
@@
(24a) $\rho_{12}\ =\$$\mathsf{L}^{T}\mathsf{L}^{*}\ \ ,$
(24b) $\rho_{24}\ =\$$\mathsf{M}\mathsf{M}^{\dagger}\ \ ,$
(24c) $\rho_{14}\ =\$$\mathsf{N}^{T}\mathsf{N}^{*}\ \ .$
For an AME state we must have $\rho_{12}=\rho_{24}=\rho_{14}=\frac{1}{4}\mathds{1}_{4}$. After taking a complex conjugate for Eqs. (24a) and (24c), these equalities imply that the matrices $\mathsf{L}$, $\mathsf{M}$, and $\mathsf{N}$, up to a factor $\frac{1}{2}$, must be unitary. Accordingly, the columns and rows of each matrix form an orthogonal basis, with all vectors normalized to $\frac{1}{2}$. Moreover, we can choose the global phase of $\mathinner{|\psi\rangle}$ so that, because of Eq. (23),
@@
$a_{0100}\ =\ a_{1000}\ =\ 0\ \ .$
Using an argument and procedure analogous to that of Proof 1: Higuchi-Sudbery in deriving Eq. (3), we obtain that both equalities are satisfied simultaneously. Assume for the moment that $a_{0011}\neq 0$ (we discuss the opposite case below). The orthogonality conditions then require the following structure of $\mathsf{L}$ (up to sign changes and column or row permutations),
@@
This proof uses the shadow inequality by Rains [Rai99, Rai00]: Denote by $\varrho_{S}=\operatorname{tr}_{S^{c}}(\varrho)$ the marginal of $\varrho$ on subsystem $S$, where $S^{c}$ denotes the complement of $S$ in $\{1,\dots,n\}$. The inequality states that for all states $\varrho$ and subsets $T\subseteq\{1,\dots n\}$, it holds that  [^2] This trace inequality has been promoted to an operator inequality in Ref. [EHGS18].
@@
The non-negativity of this expression can be understood by projecting $\varrho^{\otimes 2}$ locally onto symmetric and anti-symmetric subspaces given by $\tfrac{1}{2}(\mathds{1}+\operatorname{swap})$ and $\tfrac{1}{2}(\mathds{1}-\operatorname{swap})$ and taking the trace [Rai98]. Expanding the resulting expression using the swap trick, $\operatorname{tr}[\operatorname{swap}(A\otimes B)]=\operatorname{tr}(AB)$, yields the inequality in Eq. (32). The shadow inequality can be understood as a monogamy of entanglement constraint which governs the sharing of quantum correlations [EHGS18].
@@
Define the linear entropy $\tau_{X|\bar{X}}\ =\ 2\left(1-\operatorname{tr}(\rho_{X}^{2})\right)$, and let
@@
Gour and Wallach showed that for a four-qubit state [GBS07],
@@
where $\tau_{ABCD}=|\!\mathinner{\langle\psi|}\sigma_{y}^{\otimes 4}\mathinner{|\psi^{*}\rangle}\!|^{2}$ is non-negative due to it being an absolute value squared. However, a state with maximally mixed two-qubit marginals must have $\tau_{1}=1$ and $\tau_{2}=3/2$, which is in contradiction to inequality (36). ∎
@@
This proof follows the strategy of Ref. [MNH24, Corollary 9], which provides a complete semidefinite programming hierarchy for the existence of qubit quantum codes with given parameters ((n,K,d))(\!(n,K,d)\!). Here, $n$ denotes the number of qubits the code contains, $K$ is the dimension of the code space, and $d$ the distance. Here, we only need a relaxation of this hierarchy, which is given by the Lovász number $\vartheta$.
Let $G$ be a graph of $N$ vertices, where we write $i\sim j$ if $\{i,j\}$ is an edge in the graph. The Lovász number $\vartheta$ is given by the following semidefinite program [GL17]:
@@
Here both $a\in\mathds{R}^{N}$ and $M\in\mathds{R}^{N\times N}$ are real. It is known that the Lovász number $\vartheta$ bounds the independence number $\alpha$ of a graph, that is the maximum size of a set in which no two vertices are connected, as $\alpha(G)\leq\vartheta(G)$.
@@
Our aim is to construct a anti-commutativity graph $G$ that captures the correlations of the AME state, and to show that its theta number is too small for it to correspond to a pure state. Denote by $\sigma_{x},\sigma_{y},\sigma_{z}$ the three Pauli matrices, and by $\sigma_{0}$ the identity matrix. Let $\mathcal{P}_{n}$ be the $n$-qubit Pauli basis, that is the set of the $4^{n}$ elements of the form $E_{\alpha}=\sigma_{\alpha_{1}}\otimes\dots\otimes\sigma_{\alpha_{n}}$, where $\sigma_{\alpha_{i}}\in\{\sigma_{0},\sigma_{x},\sigma_{y},\sigma_{z}\}$. With some abuse of notation, map the multi-index $\alpha$ to an integer between 0 and $4^{n}-1$, so that $E_{0}=\mathds{1}$.
@@
where $0\leq\alpha,\beta\leq 4^{n}-1$. It can be seen that $\Gamma$ is positive semidefinite: for any $v\in\mathds{C}^{4^{n}}$ one has,
@@
due to the fact that the Pauli operators are unitary and hermitian, so that $E_{\alpha}^{2}=\mathds{1}$  [^3] Note that all Pauli operators are hermitian so we can safely drop any daggers.. Therefore the matrix $\Gamma$ satisfies: $\Gamma_{00}=1$ and $\Gamma_{\alpha 0}=\Gamma_{\alpha\alpha}$ for all $\alpha$. Furthermore, from the normalization $\operatorname{tr}(\varrho^{2})=1$ it follows that
@@
$\Gamma_{\alpha\beta}$ $=0\quad\quad\text{if}\quad 0<\operatorname{wt}(E_{\alpha}E_{\beta})<d\,$$\quad$
(42) $\Gamma_{\alpha\alpha}$ $=0\quad\quad\text{if}\quad 0<\operatorname{wt}(E_{\alpha})<d\,$,
where the weight wt of a Pauli string is the number of subsystems it acts non-trivially on. We observe that the second condition also implies that if $0<\operatorname{wt}(E_{\alpha})<d$ then also $\Gamma_{\alpha\beta}$ for all $\beta$. This is due to the fact that if a positive semidefinite matrix has a diagonal entry equal to zero, then all entries in the same row or column are zero too. Note that one could also obtain this constraint directly from the definition of $\Gamma_{\alpha\beta}$.
@@
Let the weight $\operatorname{wt}(E)$ be the number of subsystems a tensor-product operator $E$ acts non-trivially on. In light of the conditions derived in Eqs. (Proof 7: Lovász bound) and (43), let us write
@@
Then if a four-qubit AME state exists, the following semidefinite program must, due to Eq. (41), reach a value of at least $2^{4}-1$:
@@
[…]